OpenStudy (anonymous):

Richard can read 1/5 of a book in 3/5 of an hour. At this rate, how much can Richard read in one hour?

OpenStudy (anonymous):

@sammixboo @LeeEtchison

OpenStudy (anonymous):

\[ \frac{1/5}{3/5}=\frac{x}{1} \]

OpenStudy (anonymous):

We assume he will read proportionally